# Normal forms for ordinary differential operators, I

### v1

@@ Abstract
Both applications are obtained with the help of normal forms. Namely, considering the ring of ordinary differential operators $D_{1}=K[[x]][\partial]$ as a subring of a certain complete non-commutative ring $\hat{D}_{1}^{sym}$, the normal forms of differential operators mentioned here are obtained after conjugation by some invertible operator (”Schur operator”), calculated using one of the operators in a ring. Normal forms of commuting operators are polynomials with constant coefficients in the differentiation, integration and shift operators, which have a finite order in each variable, and can be effectively calculated for any given commuting operators.
@@ 1 Introduction
The first application is a new convenient explicit parametrisation of torsion free rank one sheaves on projective irreducible curves with vanishing cohomology groups. It is well known (see e.g. [4] or review [43]) that such sheaves are exactly the spectral sheaves of commutative subrings of ordinary differential operators of rank one. This result is the first step in establishing similar parametrisation for spectral sheaves of arbitrary rank, and also for spectral sheaves of commutative subrings of operators in higher dimensions (cf. [5], [41]). It is motivated by an important problem that appears in algebraic-geometric classification of commutative subrings of operators in higher dimensions – a description of torsion free sheaves with specific conditions on cohomology groups (see [41]), in particular with fixed Hilbert polynomial and some extra conditions. In the work [5] a description of Cohen-Macaulay sheaves on the spectral surface of quantum Calogero-Moser systems was given with the help of matrix problem approach due to I. Burban and Y. Drozd [2], [3] (Cohen-Macaulay sheaves form an open part of the moduli space of torsion free sheaves with fixed Hilbert polynomial, and by that reason it is important to describe them at first), however this approach meets with a difficulty to describe sheaves with specific cohomological properties (cf. [5, Sec. 6]). We expect that our new approach will help to solve this problem in an effective way in any dimension.
@@
The second application is a commutativity criterion for operators in the Weyl algebra or, more generally, in the ring of ordinary differential operators. It is motivated by the following natural question from the Burchnall-Chaundy theory. The famous Burchnall-Chaundy lemma ([6]) says that if $P,Q\in D_{1}$ are differential operators of coprime orders $p$ and $q$, then they are algebraically dependent, and satisfy a polynomial relation of the form $F(X,Y)=\alpha X^{q}\pm Y^{p}+\ldots=0$, where $GCD(p,q)=1$ and $\lambda\neq 0$ (here the weighted degree $v_{q,p}$ from [10] of $F(X,Y)$ is $pq$, and $\ldots$ mean terms of lower weighted degree, where the weight of $X$ is $p$, and the weight of $Y$ is $q$). [^2] A similar result for operators of rank $r$ was established in [38], cf. [31, Th. 2.11] Vice versa, if $P,Q\in D$ is a solution of such polynomial $f(X,Y)$ [^3] A solution of the equation $f(X,Y)=0$ is a pair $(P,Q)\in D$ such that $\sum_{i,j=0}^{n}\alpha_{ij}P^{i}Q^{j}=0$., then $[P,Q]=0$. Now a natural question whether $F(P,Q)=0\Rightarrow[P,Q]=0$ for generic polynomial $F$ appears. This question appears to be surprisingly difficult in general case. We give a partial affirmative answer on this question in the case when the normal form has the restriction top line (see discussion below).
@@
The necessity of further development of the Schur theory was not restricted only by the above mentioned applications. Let’s recall two major theorems of this theory from [42] (we formulate them here in a simplified form):
@@
If the operators $P_{i}$ from the first theorem are differential, i.e. $P_{i}\in D_{n}$, a natural question appears: what is the shape of the operator $S$? In dimension one case it is well known (see e.g. [26], [18], [34] or the book [43] and references therein) that, if the centraliser of $P_{1}$ is non-trivial, then $S$ can be expressed via the Baker-Akhieser function (and vice versa), though in quite non-trivial way, if the rank of the centraliser is greater than one. Starting from works [14], [15], [16] it is known that the Baker-Akhieser function plays an important role in many problems of mathematical physics, in particular it played a key role in constructing explicit examples of commuting operators in many works (cf. [25], [24]). Analogously, in higher dimension the Schur-Sato operator $S$ determines the common eigenfunction of commuting operators (of different nature), cf. [5, Sec. 6], [41], and the knowledge of its shape could help to prove the classification conjecture [41, Conj. 7.11] about commuting partial differential operators. Besides, in higher dimension, the operator $S$ determines an order-preserving endomorphism of the Weyl algebra ([42, Cor. 2.1]), thus giving hint to the Dixmier and Jacobian conjectures.
Analysing the shape of the Schur operator $S$, we find that all its homogeneous components are non-commutative polynomials with constant coefficients in the differentiation, integration, shift operators $A_{i}$ (see below) and the operator $\Gamma:=x\partial$ (we call such polynomials as HCPC for short), which have a finite order in each variable and additionally satisfy a specific property of being totally free of $B_{j}$ (see definition 2.6). All HCPC form a subring $Hcpc(k)$, which occasionally looks very similar to the algebra of polynomial integro-differential operators studied in the paper [1] (though the shift operators are not included into this algebra). We establish estimates on the degrees of this polynomials in section 2.4. As a result we encode all necessary properties of the operator $S$ in a condition $A_{q}(k)$ of section 2.5 (see definition 2.8 or the list of notations below). With the help of this condition, we gave a criterion of an operator $P\in\hat{D}_{1}^{sym}$ to be a differential operator (see theorem 2.2).
@@
Normal forms of a commutative subring $B\subset D_{1}$ appear to be a very effective tool to describe the moduli space of spectral sheaves, i.e. torsion free sheaves on the spectral curve with certain conditions on cohomology groups, cf. [4, §1.3] and theorem 3.1 below. Roughly speaking, the set of coefficients of a normalised normal form determines such a sheaf up to an isomorphism. This set depends on a choice of normalisation, and can be thought of as a system of local coordinates on a chart of a manifold – the moduli space of spectral sheaves. Precise statement about this parametrisation in case of sheaves of rank one is formulated in theorem 3.3.
@@
Acknowledgements. A. Zheglov is grateful to the Sino-Russian Mathematics Center at Peking University for hospitality and excellent working conditions while preparing this paper.
@@ 1.1 List of notations
  1. 1.
    $K$ is a field of characteristic zero. Recall some notation from [42]: $\hat{R}:=K[[x_{1},\ldots,x_{n}]]$, the $K$-vector space
@@
  2. 2.
    In this paper we use: $\hat{R}:=K[[x]]$, $D_{1}:=\hat{R}[\partial]$,
@@
    Operators: $\delta:=\exp((-x)\ast\partial)$, $\int:=(1-exp((-x)\ast\partial))\cdot\partial^{-1}$, $A_{k;i}:=\exp((\xi^{i}-1)x\ast\partial)\in\hat{D}^{sym}_{1}\hat{\otimes}_{K}\tilde{K}$ (in the case when $k$ is fixed, simply written as $A_{i}$), $\Gamma_{i}=(x\partial)^{i}$. $B_{n}=\frac{1}{(n-1)!}x^{n-1}\delta\partial^{n-1}$.
@@
  3. 3.
    In section 3, $\mathfrak{B}={{\cal R}}_{S}$ is the right quotient ring of ${{\cal R}}=\tilde{K}^{\oplus k}[D,\sigma]$ by $S=\{D^{k}|k\geq 0\}$. And the ring of skew pseudo-differential operators
@@
  4. 4.
    In section 4, suppose $H$ is an operator whose components are all HCP. Then $E(H)$ denotes the point set where $f_{l,i;r}\neq 0$, $v_{\sigma,\rho}$ stands for the weight degree of $H$, and $f_{\sigma,\rho}$ for the highest terms associated to $(\sigma,\rho)$:
@@ 2.1 Preliminary statements from the Schur theory for the ring $\hat{D}_{1}^{sym}$
Note that the order function $\mathop{\rm\bf ord}$ coincide with the weight function $v_{1,-1}$ on the ring $A_{1}$ from the paper [10]. For any operator $P\in D_{1}$, we define the usual order (or degree) of $P$ as $\deg(P):=v_{0,1}(P)$. If $P\in D_{1}$ has an invertible highest coefficient (with respect to the usual order), then it is easy to see that $\deg(P)=\mathop{\rm\bf ord}(P)$.
@@
Unlike the usual ring of PDOs the ring $\hat{D}_{1}^{sym}$ contains zero divisors. There are the following obvious properties of the order function (cf. the proof of [5, Th. 5.3]):
@@
The first part of lemma coincides with [42, Lem. 7.2]. The last assertion follows easily from the proof of [42, Lem. 7.2]. Namely, $A$ is a polynomial iff the infinite system of linear equations hold
@@
The following claim is a partial case of [42, Prop. 7.1].
@@
    $A_{i}A_{j}=A_{i+j}$
    $\partial^{p}A_{i}=\xi^{pi}A_{i}\partial^{p}$
    $A_{i}x^{p}=\xi^{pi}x^{p}A_{i}$
    $\int^{p}A_{i}=\xi^{-pi}A_{i}\int^{p}$
@@
For convenience of the reader we’ll give here a proof of item 2 which is easier in our case that the proof of [42, Prop. 7.1] (and we’ll use some of its arguments later).
@@
$\partial^{k}(Q_{-m})+\sum_{i=1}^{k-1}\binom{k}{i}\partial^{k-i}(Q_{-m})\partial^{i}=\\ a_{m+z}\frac{(m+z)!}{(m+z-k)!}x^{m+z-k}\partial^{z}+\sum_{p>0}a_{m+z+p}^{\prime}x^{m+z-k+p}\partial^{z+p}\neq 0$
@@
Recall that any such operator can be normalised, i.e. reduced to the form $P=\partial^{k}+c_{k-2}\partial^{k-2}+\ldots+c_{0}$, with the help of some change of variables and conjugation by invertible function, see e.g. [4, Prop. 1.3 and Rem. 1.6].
@@
All assertions of lemma can be deduced from the proof of [42, Prop.7.1]. Another proof is as follows.
@@ 7. We have
$\int(x\partial)=(\int x)\partial=(x\int-\int^{2})\partial=x(1-\delta)-\int(1-\delta)\\ =x-\int-x\delta+\int\delta=x-\int=(x\partial-1)\int.$
@@ 8. First note that
$(x\partial)B_{j}=(x\partial)\frac{1}{(j-1)!}x^{j-1}\delta\partial^{j-1}\\ =(j-1)\frac{1}{(j-1)!}x^{j-1}\delta\partial^{j-1}+\frac{1}{(j-1)!}x^{j}\partial\delta\partial^{j-1}\\ =(j-1)B_{j}+0$
@@ 2.3 Homogeneous canonical polynomials
  1. 1.
    $f_{i;r}(x,A_{k;i},\partial)$ is a polynomial of $x,A_{k;i},\partial$, $\mathop{\rm\bf ord}(f_{i;r})=0$, of the form
@@
$0=[\sum_{0\leq i<k}\sum_{0\leq l\leq d_{i}}f_{l,i;r}^{\prime}\xi^{-il}A_{k;i}\Gamma_{l}+\sum_{0<j\leq N^{\prime}}g_{j;r}^{\prime}B_{j}]B_{kn+t+1}=\\ \sum_{0\leq i<k}\sum_{0\leq l\leq d_{i}}f_{l,i;r}^{\prime}\xi^{-il}\xi^{it}(kn+t)^{l}B_{kn+t+1}$
@@
([30, Ch.2, item 10 Th.8]) Suppose $L,f\in\tilde{K}[t]$ are non-zero polynomials with $\deg f=r$, $\lambda\in\tilde{K}$, and $q$ is the multiplicity of $t-\lambda$ in the polynomial $L(t)$ (if $L(\lambda)\neq 0$, then $q=0$). Then the ODE
@@
  1. 1.
    We use the same idea as in the proof of proposition 2.1. Considering $\partial$ as a constant the original equation becomes an ODE in the ring $\tilde{R}$:
@@
  2. 2.
    We have two possible case of such monomials:
    1. (a)

### v3

@@ Abstract
Both applications are obtained with the help of normal forms. Namely, considering the ring of ordinary differential operators $D_{1}=K[[x]][\partial]$ as a subring of a certain complete non-commutative ring $\hat{D}_{1}^{sym}$, the normal forms of differential operators mentioned here are obtained after conjugation by some invertible operator (”Schur operator”), calculated using one of the operators in a ring. Normal forms of commuting operators are polynomials with constant coefficients in the differentiation, integration and shift operators, which have a restricted finite order in each variable, and can be effectively calculated for any given commuting operators.
@@ 1 Introduction
The first application is a new convenient explicit parametrisation of torsion free rank one sheaves on projective irreducible curves with vanishing cohomology groups. It is well known (see e.g. [5] or review [47]) that such sheaves are exactly the spectral sheaves of commutative subrings of ordinary differential operators of rank one. This result is the first step in establishing similar parametrisation for spectral sheaves of arbitrary rank, and also for spectral sheaves of commutative subrings of operators in higher dimensions (cf. [6], [45]). It is motivated by an important problem that appears in algebraic-geometric classification of commutative subrings of operators in higher dimensions – a description of torsion free sheaves with specific conditions on cohomology groups (see [45]), in particular with fixed Hilbert polynomial and some extra conditions. In the work [6] a description of Cohen-Macaulay sheaves on the spectral surface of quantum Calogero-Moser systems was given with the help of matrix problem approach due to I. Burban and Y. Drozd [3], [4] (Cohen-Macaulay sheaves form an open part of the moduli space of torsion free sheaves with fixed Hilbert polynomial, and by that reason it is important to describe them at first), however this approach meets with a difficulty to describe sheaves with specific cohomological properties (cf. [6, Sec. 6]). We expect that our new approach will help to solve this problem in an effective way in any dimension.
The second application is a commutativity criterion for operators in the Weyl algebra or, more generally, in the ring of ordinary differential operators. It is motivated by the following natural question from the Burchnall-Chaundy theory.
@@
The famous Burchnall-Chaundy lemma ([7]) says that any two commuting differential operators $P,Q\in D_{1}:=K[[x]][\partial]$ are algebraically dependent. More precisely, if the orders $n,m$ of operators $P,Q$ are coprime111i.e. the rank of the ring $K[P,Q]$ is 1, see e.g. [47] for relevant definitions, in particular [47, Lemma 5.23] for a proof of the Burchnall-Chaundy lemma in general form. The statement about the form of polynomial follows easily from the proof., then there exists an irreducible polynomial $f(X,Y)$ of weighted degree $v_{n,m}(f)=mn$ of special form (here the weighted degree is defined as in Dixmier’s paper [14], cf. item 4 in the List of Notations below): $f(X,Y)=\alpha X^{m}\pm Y^{n}+\ldots$ (here $\ldots$ mean terms of lower weighted degree, $0\neq\alpha\in K$; in particular, for coprime $p$ and $q$ the polynomial $f$ is automatically irreducible), such that $f(P,Q)=0$. A similar result for commuting operators of rank $r$ was established in [42] (cf. [35, Th. 2.11]), in this case $m=\mathop{\rm ord}\nolimits(Q)/r$, $n=\mathop{\rm ord}\nolimits(P)/r$, and again $GCD(m,n)=1$. Vice versa, if $P,Q\in D_{1}$ is a solution of such polynomial $f(X,Y)$ [^2] A solution of the equation $f(X,Y)=0$ is a pair $(P,Q)\in D$ such that $\sum_{i,j=0}^{n}\alpha_{ij}P^{i}Q^{j}=0$., then $[P,Q]=0$. Now a natural question whether $F(P,Q)=0\Rightarrow[P,Q]=0$ for generic polynomial $F$ appears. This question appears to be surprisingly difficult in general case. We give a partial affirmative answer on this question in the case when the normal form has the restriction top line (see discussion below).
@@
The necessity of further development of the Schur theory was not restricted only by the above mentioned applications. Let’s recall two major theorems of this theory from [46] (we formulate them here in a simplified form, for notations see the List of notations below):
@@
If the operators $P_{i}$ from the first theorem are differential, i.e. $P_{i}\in D_{n}$, a natural question appears: what is the shape of the operator $S$? In dimension one case it is well known (see e.g. [30], [22], [38] or the book [47] and references therein) that, if the centraliser of $P_{1}$ is non-trivial, then $S$ can be expressed via the Baker-Akhieser function (and vice versa), though in quite non-trivial way, if the rank of the centraliser is greater than one. Starting from works [18], [19], [20] it is known that the Baker-Akhieser function plays an important role in many problems of mathematical physics, in particular it played a key role in constructing explicit examples of commuting operators in many works (cf. [29], [28] for examples in dimension one, and [10], [11], [12] for examples in higher dimensions). Analogously, in higher dimension the Schur-Sato operator $S$ determines the common eigenfunction of commuting operators (of different nature), see [6, Sec. 6], [45], cf. also [2] and references therein, and the knowledge of its shape could help to prove the classification conjecture [45, Conj. 7.11] about commuting partial differential operators. Besides, in higher dimension, the operator $S$ determines an order-preserving endomorphism of the Weyl algebra ([46, Cor. 2.1]), thus giving hint to the Dixmier and Jacobian conjectures.
Analysing the shape of the Schur operator $S$, we find that all its homogeneous components are non-commutative polynomials with constant coefficients in the differentiation, integration, shift operators $A_{i}$ (see below) and the operator $\Gamma:=x\partial$ (we call such polynomials as HCPC for short), which have a finite order in each variable and additionally satisfy a specific property of being totally free of $B_{j}$ (see definition 2.6). All HCPC form a subring $Hcpc(k)$, which occasionally looks very similar to the algebra of polynomial integro-differential operators studied in the paper [1] (though the shift operators are not included into this algebra). We establish estimates on the degrees of these polynomials in section 2.4. As a result we encode all necessary properties of the operator $S$ in a condition $A_{q}(k)$ of section 2.5 (see definition 2.8 or the list of notations below). With the help of this condition, we gave a criterion of an operator $P\in\hat{D}_{1}^{sym}$ to be a differential operator (see theorem 2.2).
@@
Normal forms of a commutative subring $B\subset D_{1}$ appear to be a very effective tool to describe the moduli space of spectral sheaves, i.e. torsion free sheaves on the spectral curve with certain conditions on cohomology groups, cf. [5, §1.3] and theorem 3.1 below. Roughly speaking, the set of coefficients of a normalised normal form determines such a sheaf up to an isomorphism. This set depends on a choice of normalisation, and can be thought of as a system of local coordinates on a chart of a manifold – the moduli space of spectral sheaves. Precise statement about this parametrisation in case of sheaves of rank one is formulated in theorem 3.3.
@@
Acknowledgements. This work of the first author was partially supported by the National Key R and D Program of China (Grant No. 2020YFE0204200). The work of the second author was partially supported by RSF grant no. 22-11-00272.
A. Zheglov is grateful to the Sino-Russian Mathematics Center at Peking University for hospitality and excellent working conditions while preparing this paper.
We are grateful to Huijun Fan for his interest in our work, and for his advice and support to J. Guo.
@@ 1.1 List of notations
1. $K$ is a field of characteristic zero. Recall some notation from [46]: $\hat{R}:=K[[x_{1},\ldots,x_{n}]]$, the $K$-vector space
@@
2. In this paper we use: $\hat{R}:=K[[x]]$, $D_{1}:=\hat{R}[\partial]$,
@@
Operators: $\delta:=\exp((-x)\ast\partial)$, $\int:=(1-exp((-x)\ast\partial))\cdot\partial^{-1}$, $A_{k;i}:=\exp((\xi^{i}-1)x\ast\partial)\in\hat{D}^{sym}_{1}\hat{\otimes}_{K}\tilde{K}$ (in the case when $k$ is fixed, simply written as $A_{i}$), where $\tilde{K}=K[\xi]$, $\xi$ is a primitive $k$th root of unity, $\Gamma_{i}=(x\partial)^{i}$. $B_{n}=\frac{1}{(n-1)!}x^{n-1}\delta\partial^{n-1}$.
The operator $P\in D_{1}$ is called normalized if $P=\partial^{p}+a_{p-2}\partial^{p-2}+\ldots$. The operator $P\in D_{1}$ is monic if its highest coefficient is 1. Analogously, $P\in\hat{D}_{1}^{sym}$ is monic if $\sigma(P)=\partial^{p}$.
@@
3. In section 3, $\mathfrak{B}={{\cal R}}_{S}$ is the right quotient ring of ${{\cal R}}=\tilde{K}^{\oplus k}[D,\sigma]$ by $S=\{D^{k}|k\geq 0\}$. And the ring of skew pseudo-differential operators
@@
is an isomorphism of $\tilde{K}$-algebras.
The map
$\hat{\Phi}:\widehat{Hcpc}_{B}(k)\hookrightarrow E_{k}$
denotes an embedding of $\tilde{K}$-algebras (see lemma 3.2).
@@
4. In section 4, suppose $H$ is an operator whose components are all HCP. Then $E(H)$ denotes the point set where $f_{l,i;r}\neq 0$, $v_{\sigma,\rho}$ stands for the weight degree of $H$, and $f_{\sigma,\rho}$ for the highest terms associated to $(\sigma,\rho)$:
@@ 2.1 Preliminary statements from the Schur theory for the ring $\hat{D}_{1}^{sym}$
Note that the order function $\mathop{\rm\bf ord}$ coincide with the weight function $v_{1,-1}$ on the ring $A_{1}$ from the paper [14]. For any operator $P\in D_{1}$, we define the usual order (or degree) of $P$ as $\deg(P):=v_{0,1}(P)$. If $P\in D_{1}$ has an invertible highest coefficient (with respect to the usual order), then it is easy to see that $\deg(P)=\mathop{\rm\bf ord}(P)$. Vice versa, any $P\in D_{1}$ with $\deg(P)=\mathop{\rm\bf ord}(P)$ has an invertible highest coefficient.
@@
Unlike the usual ring of PDOs the ring $\hat{D}_{1}^{sym}$ contains zero divisors. There are the following obvious properties of the order function (cf. the proof of [6, Th. 5.3]):
@@
The first part of lemma coincides with [46, Lem. 7.2]. The last assertion follows easily from the proof of [46, Lem. 7.2]. Namely, $A$ is a polynomial iff the infinite system of linear equations hold
@@
The following claim is a partial case of [46, Prop. 7.1].
@@
    $A_{i}A_{j}=A_{i+j}\quad\partial^{p}A_{i}=\xi^{pi}A_{i}\partial^{p}\quad A_{i}x^{p}=\xi^{pi}x^{p}A_{i}\quad\int^{p}A_{i}=\xi^{-pi}A_{i}\int^{p}.$
@@
For convenience of the reader we’ll give here a proof of item 2 which is easier in our case that the proof of [46, Prop. 7.1] (and we’ll use some of its arguments later).
@@
$\partial^{k}(Q_{-m})+\sum_{i=1}^{k-1}\binom{k}{i}\partial^{k-i}(Q_{-m})\partial^{i}=a_{m+z}\frac{(m+z)!}{(m+z-k)!}x^{m+z-k}\partial^{z}+\sum_{p>0}a_{m+z+p}^{\prime}x^{m+z-k+p}\partial^{z+p}\neq 0$
@@
Recall that any such operator can be normalised, i.e. reduced to the form $P=\partial^{k}+c_{k-2}\partial^{k-2}+\ldots+c_{0}$, with the help of some change of variables and conjugation by invertible function, see e.g. [5, Prop. 1.3 and Rem. 1.6].
@@
All assertions of lemma can be deduced from the proof of [46, Prop.7.1]. Another proof is as follows.
@@ 7. We have
$\int(x\partial)=(\int x)\partial=(x\int-\int^{2})\partial=x(1-\delta)-\int(1-\delta)=x-\int-x\delta+\int\delta=x-\int=(x\partial-1)\int.$
@@ 8. First note that
$(x\partial)B_{j}=(x\partial)\frac{1}{(j-1)!}x^{j-1}\delta\partial^{j-1}=(j-1)\frac{1}{(j-1)!}x^{j-1}\delta\partial^{j-1}+\frac{1}{(j-1)!}x^{j}\partial\delta\partial^{j-1}=(j-1)B_{j}+0$
@@ 2.3 Homogeneous canonical polynomials
1. $f_{i;r}(x,A_{k;i},\partial)$ is a polynomial of $x,A_{k;i},\partial$, $\mathop{\rm\bf ord}(f_{i;r})=0$, of the form
@@
$0=[\sum_{0\leq i<k}\sum_{0\leq l\leq d_{i}}f_{l,i;r}^{\prime}\xi^{-il}A_{k;i}\Gamma_{l}+\sum_{0<j\leq N^{\prime}}g_{j;r}^{\prime}B_{j}]B_{kn+t+1}=\sum_{0\leq i<k}\sum_{0\leq l\leq d_{i}}f_{l,i;r}^{\prime}\xi^{-il}\xi^{it}(kn+t)^{l}B_{kn+t+1}$
@@
([34, Ch.2, item 10 Th.8]) Suppose $L,f\in\tilde{K}[t]$ are non-zero polynomials with $\deg f=r$, $\lambda\in\tilde{K}$, and $q$ is the multiplicity of $t-\lambda$ in the polynomial $L(t)$ (if $L(\lambda)\neq 0$, then $q=0$). Then the ODE
@@
1. We use the same idea as in the proof of proposition 2.1. Considering $\partial$ as a constant the original equation becomes an ODE in the ring $\tilde{R}$:
@@
2. We have two possible case of such monomials:
  1. 1.